A person is driving at an average speed of 60 [mi/h] for 23 minutes.

a) How far does the person travel within this time period?

b) Convert this distance to miles and show your work.


1
Expert's answer
2021-09-20T10:00:04-0400

Given:

"v=60\:\rm mi/hr=26.82\: m/s"

"t=23\:\rm min=1380\: s"


The distance traveled by a person

(a)

"d=vt=26.82*1380=37014 \:\rm m"

(b)

"37014 \:\rm m=37014 *\rm \frac{1\: m}{1609\: m/mi}=23\: mi"